What is the pH of a buffer solution containing 0.63 NaHCO3 and 0.32 H2CO3?


1
Expert's answer
2021-10-05T01:43:52-0400

pH = pKa+log(acid/base)

pKa= 3.75

Log(0.63/0.32)=0.294

= 3.75+0.294

pH= 4.04